# Hybrid discrimination strategy based on photon-number-resolving detectors and mesoscopic twin-beam states

### v1

@@ II.1 Thermal states and mesoscopic TWB states
where $\mu$ is the effective number of modes, while $\langle m\rangle$ is the mean number of detected photons. The photon-number statistics in each arm is super-Poissonian and can be also discriminated by means of the Fano factor. The latter is defined in terms of detected photons as $F(m)=\sigma^{2}(m)/\langle m\rangle=\eta F(n)+(1-\eta)$, where $\sigma^{2}(m)$ is the variance of the distribution of the number of detected photons and $\eta$ is the quantum efficiency of the detector [20]. It is well known that $F=1$ in the case of Poissonian light, $F<1$ for sub-Poissonian light (nonclassical), and $F>1$ for super-Poissonian light.
The two noise signals are sent in an alternative way through a communication channel together with a portion of a multi-mode TWB state. This is a quantum-correlated optical state that can be expressed in terms of the following density matrix [21]
$\rho^{\mu}_{\rm TWB}=\sum_{n=0}^{\infty}P_{\rm mth}(n)|n,n\rangle\langle n,n|,$ (2)
in which
$|n\rangle=\delta_{n,\,(\sum_{k=1}^{\mu}n_{k})}\bigotimes_{k=1}^{\mu}|n_{k}\rangle_{k}$ (3)
and $n$ is the overall number of photons, while $P_{\rm mth}(n)$ is the multi-mode thermal photon-number distribution in Eq. (1) written in terms of incident photons. This means that the measurements performed by Bob return a mean value that is given by the sum of the mean value of the thermal signal and that of the TWB portion. The other TWB portion is retained and measured by Alice herself. The presence of the TWB in the communication channel allows for revealing possible eavesdropping attacks. For instance, this can be achieved by applying nonclassicality criteria involving the number of photons, such as that based on the noise reduction factor, $R$. In terms of the number of detected photons, it reads as [22]
@@
where $\langle m\rangle$ is the mean number of detected photons in a TWB arm, $\eta$ is the global quantum efficiency of the detection system, $t\in(0,1)$ is the transmission efficiency quantifying the balancing between the two arms, while $\langle m_{\rm N}\rangle$ and $\mu_{\rm N}$ are the mean value and the effective number of modes of the thermal noise signal, respectively. This means that, in general, the value of $R$ increases in the presence of noise, since the latter is not correlated with the TWB, and also in the presence of losses, as $t<1$. In this scenario, nonclassical correlations, $i.e.$ $R<1$, can be detected provided that
@@
Figure 1: Sketch of the experimental setup (a) for the characterization of TWB states and (b) for the communication protocol. HWP: half-wave plate; BBO: $\beta$-barium-borate crystal; BF: band-pass filter; PH: pin-hole; AD: achromatic doublet; MF: multi-mode optic fiber; SiPM: silicon photomultiplier; M: mirror; PBS: polarizing beam splitter; CL: convex lens; ND: neutral density filter; GGD: ground-glass disk. The stage of beam preparation comprises two prisms and a demagnifying telescope. The pump field is indicated in purple, the signal (idler) beam is shown as the upper (lower) red line, and in (b) the noise signal, generated by a different laser source, is indicated in green. The dashed line denotes the noise signal is superimposed on the signal arm of the TWB in post-processing by suitably combining the acquired data.
@@ III.1 Communication channel
Outside the BBO, spatial and spectral filtering (identical in the two arms) allows to select two entangled regions. As to the spectral filtering, we use two band-pass filters, having a central wavelength of 692 nm on one arm and 697 nm on the other arm, and both having a width of $\sim 20$ nm. For the spatial filtering, we use an adjustable pin-hole having a diameter of 1 cm. The filtered light in each arm is then focused by an achromatic doublet into a multi-mode optical fiber having a 1-mm-core diameter, and then delivered to the detector. The detection is performed by means of two SiPMs (model S13360-3050CS, Hamamatsu Photonics) [30]. These are solid-state PNR detectors endowed with hundreds of cells (pixels), operated in the Geiger-Müller regime. Under the assumption that each cell is fired by at most one photon impinging on it, each SiPM yields a single output that is proportional to the number of fired cells. The detectors are plugged into a power supply and amplification unit (PSAU, model SP5600, CAEN) that supplies a bias voltage of 56 V to the SiPMs and amplifies the output by 17 dB. Each output is synchronously integrated over an interval of 10 ns by a boxcar-gated integrator and acquired.
@@ III.2 Binary noise signal
The experimental setup for the study of the discrimination protocol is sketched in Fig. 1(b). It is equivalent to the previous one [Fig. 1(a)] except for the additional optical elements required to generate the noise signal and to superimpose it on the signal arm of the TWB. The thermal noise is generated by sending picosecond pulses at 523 nm to a ground-glass disk. A single speckle is selected in far field by an iris 300-$\mu$m large. The energy of the laser beam is changed by means of a variable neutral density filter. In this way, it is possible to produce the two noise signals with different mean values, necessary to encode the logical bits 0 and 1, that can be superimposed in an alternative way to the portion of TWB sent to Bob. In the present study, we superimpose the noise signals in post-processing [see the dashed line in Fig. 1(b)], following a similar procedure already adopted in [31, 32] to merge the data corresponding to TWB states and those corresponding to noise signals. The practical implementation of the communication channel based on TWB is obtained by inserting some optical elements that increase its length and introduce some losses. Despite the latter, nonclassical correlations can still be detected by Bob, thus enabling the devised communication protocol.
@@ IV.1 Measurements and data processing
The quantities we are interested in—the mean number of detected photons $\langle m\rangle$ and the noise reduction factor $R$—ultimately depend on the intensity of the light that impinges on the two detectors. Therefore, it is necessary to preliminarily characterize the TWB states for different values of their intensity in the mesoscopic intensity domain, and the two noise signals we superimpose on the signal arm to encode the bits. Concerning the first aspect, we use a combination of HWP and PBS [Fig. 1(a)] as a knob to tune the intensity of the pump beam and so, ultimately, to change the mean value of TWB. In more detail, we consider 4 angle values of the HWP placed on the pump beam path to investigate the statistical properties of the TWB alone as a function of the pump beam intensity. The TWB is then investigated in the presence of losses by using the asymmetric setup in Fig. 1(b). The two noise signals are also preliminarily characterized in order to verify that we have at our disposal single-mode thermal states. Their mean values are properly selected in order to satisfy Eq. (6), thus ensuring $R<1$.
@@
Figure 3: Characterization of the symmetric TWB. (a) Detected-photon distribution $P(m_{k})$ in the idler $k=i$ (blue-scale-colored diamonds) and signal $k=s$ (red-scale-colored circles). The fitting function (line) is the multi-mode thermal distribution (MMTD) in Eq. (1). Light (dark) colors correspond to low (high) $\langle m_{k}\rangle$. (b) Fano factor $F(m_{k})>1$ of each arm as a function of $\langle m_{k}\rangle$. (c) Noise reduction factor $R<1$ as a function of $\langle m_{s}\rangle$. Results in (b) and (c) are obtained by dividing the $10^{5}$ data of each dataset into 4 batches.
@@ IV.2 Characterization of the communication channel
As explained in Sec. II, many properties of the TWB that qualify the communication channel can be determined starting from the number of detected photons $m_{k}$ in each arm, signal ($k=s$) and idler ($k=i$). Among them, here we consider the photon-number distribution, the Fano factor, and the noise reduction factor. For all the four increasing values of the pump intensity, our results are consistent with the expected behavior of the TWB. The two arms are well balanced, $\langle m_{s}\rangle\approx\langle m_{i}\rangle$, and each detected-photon distribution is well fitted by the multi-mode thermal distribution in Eq. (1), as shown in Fig. 3(a). The super-Poissonian nature of the measured light is also evidenced by the Fano factor, which is larger than 1 for all the datasets shown in Fig. 3(b). Finally, the noise reduction factor is $R<1$ [see Fig. 3(c)] in all cases, denoting the presence of nonclassical photon-number correlations in the mesoscopic intensity domain. By fitting the latter results using Eq. (5) with $t=1$ (lossless configuration, the two arms are balanced) and $\langle m_{\rm N}\rangle=0$ (noiseless configuration), we can estimate the detection efficiency $\eta=(8.5\pm 0.7)\%$, which is lower than the expected one ($\eta<15\%$ at 694 nm, according to the datasheet[30]). The discrepancy between our estimate and the expected upper bound to $\eta$ is likely due to possible losses in the collection of light. Remarkably, despite its low value, this efficiency is still sufficient to detect nonclassical correlations.
Figure 4: Characterization of the asymmetric TWB and the noise signals. (a) Detected-photon distribution $P(m_{k})$ in the idler $k=i$ and signal $k=s$ and (b) in the two noise signals (NS). The fitting function (line) is the multi-mode thermal distribution (MMTD) in Eq. (1). (c) Noise reduction factor of the TWB with superimposed noise signal as a function of $\langle m_{\rm N}\rangle$. The dashed line indicates the value of $R$ with associated error in the noiseless lossy configuration. Results in (c) are obtained by dividing the $10^{5}$ data of each dataset into 4 batches.
@@
Figure 5: Effects of the batch size on the estimation of $R$ and $\langle m_{s}\rangle$. Mean value and associated error of the noise reduction factor $R$ (a) and of the mean number of detected photons in the signal arm $\langle m_{s}\rangle$ (b) as a function of the batch size $d_{b}$. Histogram of the values of $R$ (c) and of $\langle m_{s}\rangle$ (d) for $d_{b}=20000$ (top) and $d_{b}=40000$ (bottom). In (c,d) the vertical dashed line denotes the discrimination threshold $R_{th}=0.968$ and $\langle m_{s}\rangle_{th}=3.72$, respectively. Results are obtained from $5000$ batches drawn from the original datasets by bootstrap procedure.
@@ IV.3 Performance of the discrimination strategy
Figure 6: Effects of the batch size and threshold on the discrimination of bits. (a) Error probability (misclassification) as a function of the batch size $d_{b}$ when using $R$ (blue diamonds) or $\langle m_{s}\rangle$ (red circles) to discriminate bits 0 or 1 for a given threshold (see Fig. 5). (b) ROC of each discrimination strategy: True positive rate (TPR) as a function of the false positive rate (FPR) for two representative batch sizes $d_{b}=20000,40000$. The perfect classifier corresponds to (FPR=0,TPR=1) (yellow shadow), while the random classifier to the line TPR=FPR (grey dotted line). The empty markers denote in (a) data associated with $d_{b}=20000,40000$, and in (b) those associated with threshold values $R_{th}=0.968$ and $\langle m_{s}\rangle_{th}=3.72$. Results in (a)-(b) are obtained from $5000$ batches drawn from the original dataset by bootstrap procedure.
@@ V Proof-of-principle test
Figure 7: Example of discrimination of a 400-bit key for two batch sizes $d_{b}=20000,40000$. The raw key is visualized as a 20×20 binary image, where each pixel corresponds to a single bit. Alice generates her key using a pseudo-random number generator, while Bob obtains his key by measuring either $\langle m_{s}\rangle$ or $R$ and discriminating bits according to the thresholds reported in Fig. 6. Errors in Bob’s keys, highlighted in light- and dark-red, occur with rates determined from $5000$ batches drawn from the original dataset by bootstrap procedure.
As expected, Bob can exactly decode Alice’s raw key using the $\langle m_{s}\rangle$-based discrimination strategy, while the raw key he obtains using the $R$-based strategy is affected by errors. In particular, the error probability decreases from $P_{\hbox{\tiny$R$}}^{\rm(err)}\approx 24\%$ to $\approx 16\%$ upon increasing the batch size from $d_{b}=20000$ to $40000$ [see also Fig. 6(a)]. In detail, the false positive/negative rates decreases from ${\rm FPR}\approx 19\%$ to $\approx 11\%$ and ${\rm FNR}\approx 29\%$ to $\approx 22\%$. Therefore, a bit-1 decoded as a bit-0 (i.e., a false negative) constitutes a misclassification error that occurs systematically more often than the opposite case, where a bit-0 is decoded as a bit-1 (i.e., a false positive). This asymmetry in error rates arises from the choice of the discrimination threshold $R_{th}$ in the present example. According to the analysis of the ROC curve, and recalling that ${\rm FNR}\equiv 1-{\rm TPR}$, the condition ${\rm FNR}={\rm FPR}$ corresponds to points lying along the anti-diagonal ${\rm TPR}=1-{\rm FPR}$. However, as illustrated in Fig. 6(b), the points associated with $R_{th}=0.968$ lie below the anti-diagonal, thus indicating ${\rm FNR}>{\rm FPR}$ (see also Fig. 5(c), which shows that the threshold $R_{th}$ lies closer to one distribution than to the other).
Since this protocol requires estimating $R$, each bit must be encoded using a batch size at least $d_{b}\sim O(10^{4})$. In this regard, $d_{b}\approx 40000$ represents a suitable batch size to reliably estimate $R$ while keeping the number of shots per bit reasonably limited.
@@
Figure 8: An ideal intercept-resend attack can be revealed by the values of the noise reduction factor $R$. Eve intercepts $d_{\rm E}$ shots of a batch encoding a bit ($d_{b}=40000$) and resends, in place of the intercepted light, $d_{\rm E}$ shots belonging to a Poissonian light having the same mean value as that of the intercepted light. While the mean value measured by Bob is not affected by this attack (data not shown), the value of the noise reduction factor is. Blue color refers to bit 0, while red color refers to bit 1. The blue dotted line + blue highlighted band corresponds to $R^{(0)}\pm 2\sigma$, while the red dashed line + red highlighted band corresponds to $R^{(1)}\pm 2\sigma$. The values of $R$ corresponding to the intercepted-resent bit 0 (1) are shown as blue open squares (full red squares). Results are averaged over $5000$ realizations of the intercept-resend attack.
To better investigate the feasibility of the strategy, we simulate an eavesdropping attack by assuming that Eve intercepts part of a batch encoding either 0 or 1, and re-sends a signal with the same mean value as the stolen portion of light. In doing so, Eve replaces a portion of the shots in the signal arm (Bob)—originally nonclassically correlated with those in the idler arm (Alice)—with others from a different light source, which is uncorrelated with Alice’s. This degradation of nonclassical correlations due to the intercept-resend attack is therefore expected to increase the noise reduction factor [see Eq. (5)]. In particular, we consider a synthetic signal generated according to a Poissonian distribution, since this is the statistical distribution that better approximates the distribution of the signal received by Bob in the absence of eavesdropping. The results of this intercept-resend attack are shown in Fig. 8, where the noise reduction factor is shown as a function of the ratio $d_{\rm E}/d_{b}$ between the batch size changed by Eve and that encoding either 0 or 1. The linear dependence of $R$ on $d_{\rm E}/d_{b}$ can be proved analogously to the derivation of Eq. (5). The highlighted bands indicate the uncertainty corresponding to ±$\pm 2\sigma$. It can be noticed that for bit 0 the value of $R$ exceeds $R^{(0)}+2\sigma$ and $1$ at $d_{\rm E}/d_{b}\approx 9\%$ and $\approx 84\%$, respectively. Instead, for bit 1 it exceeds $R^{(1)}+2\sigma$ and $1$ at $d_{\rm E}/d_{b}\approx 14\%$ and $\approx 78\%$, respectively. This means that Bob can easily detect the presence of Eve by simply checking whether the values of $R$ are larger than $2\sigma$. On the contrary, values $R>1$ are more difficult to observe, at least when Eve, as in the present example, is equipped with an ideal detector that allows perfect estimation of the mean value of the intercepted signal. At the same time, this proof-of-principle test does not take into account possible optimization strategies, such as practical reconciliation or privacy amplification, devised by Bob to improve its receiver.

### v2

# Hybrid discrimination strategy in quantum communication based on photon-number-resolving detectors and mesoscopic twin-beam states

@@ II.1 Thermal states and mesoscopic TWB states
where $\mu$ is the effective number of modes, while $\langle m\rangle$ is the mean number of detected photons. The photon-number statistics is super-Poissonian and can be also discriminated by means of the Fano factor. The latter is defined in terms of detected photons as $F(m)=\sigma^{2}(m)/\langle m\rangle=\eta F(n)+(1-\eta)$, where $\sigma^{2}(m)$ is the variance of the distribution of the number of detected photons and $\eta$ is the quantum efficiency of the detector [20]. It is well known that $F=1$ in the case of Poissonian light, $F<1$ for sub-Poissonian light (nonclassical), and $F>1$ for super-Poissonian light.
The two noise signals are sent in an alternative way through a communication channel together with a portion of a multi-mode TWB state [21], which is a quantum-correlated optical state. If the $\mu$ modes of each of the two parties of the TWB are equally populated, then the multi-mode TWB state is ideally described by a tensor product of $\mu$ identical TWB states [22, 23],
$|\Psi^{\mu}_{\rm TWB}\rangle=\bigotimes_{k=1}^{\mu}\sqrt{1-\lambda^{2}}\sum_{\nu=0}^{\infty}\lambda^{\nu}|\nu_{k}\rangle\otimes|\nu_{k}\rangle,$ (2)
where $k$ labels the modes, $\nu$ is the photon number in each mode, $\lambda$ is defined through $\lambda^{2}=\langle n\rangle/(\mu+\langle n\rangle)$, and $\langle n\rangle$ is the mean total number of photons in either of the two parties of the TWB, with $n=\sum_{k=1}^{\mu}\nu_{k}$. The state in Eq. (2) is characterized by perfect photon-number correlations between the two parties of the TWB in each mode $k$. It is possible to prove that the probability of observing an overall number of photons $n$ in each arm of the TWB is given by the multi-mode thermal photon-number distribution in Eq. (1) written in terms of the number $n$ of incident photons [24, 25, 21]. The measurements performed by Bob return a mean value that is given by the sum of the mean value of the thermal signal and that of the TWB portion. The other TWB portion is retained and measured by Alice herself. The presence of the TWB in the communication channel allows for revealing possible eavesdropping attacks. For instance, this can be achieved by applying nonclassicality criteria involving the number of photons, such as that based on the noise reduction factor, $R$. In terms of the number of detected photons, it reads as [26]
@@
where $\langle m\rangle$ is the mean number of detected photons in a TWB arm, $\eta$ is the global quantum efficiency of the detection system (including the detection efficiency and the coupling of light into fibers), $t\in[0,1]$ is the transmission efficiency quantifying the balancing between the two arms, while $\langle m_{\rm N}\rangle$ and $\mu_{\rm N}$ are the mean value and the effective number of modes of the thermal noise signal, respectively. This means that, in general, the value of $R$ increases in the presence of noise, since the latter is not correlated with the TWB, and also in the presence of losses, as $t<1$. In this scenario, nonclassical correlations, $i.e.$ $R<1$, can be detected provided that
@@
Physically, this inequality can be interpreted as follows: in the case of perfect balance between the TWB arms (i.e., for $t=1$), the nonclassicality condition requires that the mean value of the signal noise, $\langle m_{\rm N}\rangle$, be less than a term proportional to $\sqrt{2\langle m\rangle}$, where $2\langle m\rangle$ is the shot-noise level, corresponding to the total mean number of detected photons in the two TWB arms; for $t<1$, the presence of an imbalance imposes a more stringent constraint on the thermal noise signal, requiring $\langle m_{\rm N}\rangle$ to be lower than that in the balanced case.
@@ II.2 Discrimination of the noise signal as a binary classification problem
Figure 1: Sketch of the experimental setup (a) for the characterization of TWB states and (b) for the communication protocol. HWP: half-wave plate; BBO: $\beta$-barium-borate crystal; BF: band-pass filter; PH: pin-hole; AD: achromatic doublet; MF: multi-mode optic fiber; SiPM: silicon photomultiplier; M: mirror; PBS: polarizing beam splitter; CL: convex lens; ND: neutral density filter; GGD: ground-glass disk. The stage of beam preparation (dashed box) comprises two prisms and a demagnifying telescope, as detailed in [31]. The pump field is indicated in purple, the signal (idler) beam is shown as the upper (lower) red line, and in (b) the noise signal, generated by a different laser source, is indicated in green. The dashed line denotes that the noise signal is superimposed on the signal arm of the TWB in post-processing by suitably combining the acquired data.
@@ III.1 Communication channel
Outside the BBO, spatial and spectral filtering (identical in the two arms) allows to select two entangled regions. As to the spectral filtering, we use two band-pass filters, having a central wavelength of 692 nm on one arm and 697 nm on the other arm, and both having a width of $\sim 20$ nm. For the spatial filtering, we use an adjustable pin-hole having a diameter of 1 cm. The filtered light in each arm is then focused by an achromatic doublet into a multi-mode optical fiber having a 1-mm-core diameter, and then delivered to the detector. The detection is performed by means of two SiPMs (model S13360-3050CS, Hamamatsu Photonics) [34]. These are solid-state PNR detectors endowed with hundreds of cells (pixels), operated in the Geiger-Müller regime. Under the assumption that each cell is fired by at most one photon impinging on it, each SiPM yields a single output that is proportional to the number of fired cells. The SiPM model used in this work is characterized by a quantum efficiency $\sim 16\%$ at 693 nm, which can reach higher values, $\sim 40\%$, in the blue–green spectral region [31]. The detectors are plugged into a power supply and amplification unit (PSAU, model SP5600, CAEN) that supplies a bias voltage of 56 V to the SiPMs and amplifies the output by 17 dB. Each output is synchronously integrated over an interval of 10 ns by a boxcar-gated integrator and acquired.
@@ III.2 Binary noise signal
[…]
@@ IV.1 Measurements and data processing
The quantities we are interested in—the mean number of detected photons $\langle m\rangle$ and the noise reduction factor RR—ultimately depend on the intensity of the light that impinges on the two detectors. Therefore, it is necessary to preliminarily characterize the TWB states for different values of their intensity in the mesoscopic intensity domain, and the two noise signals we superimpose on the signal arm to encode the bits. Concerning the first aspect, we use a combination of HWP and PBS [Fig. 1(a)] as a knob to tune the intensity of the pump beam and so, ultimately, to change the mean value of TWB. In more detail, we consider 4 angle values of the HWP placed on the pump beam path to investigate the statistical properties of the TWB alone as a function of the pump beam intensity. The TWB is then investigated in the presence of losses by using the asymmetric setup in Fig. 1(b). The two noise signals are also preliminarily characterized in order to verify that we have at our disposal single-mode thermal states. Their mean values are properly selected in order to satisfy Eq. (5), thus ensuring $R<1$.
@@ IV.2 Characterization of the communication channel
As explained in Sec. II, many properties of the TWB that qualify the communication channel can be determined starting from the number of detected photons $m_{k}$ in each arm, signal ($k=s$) and idler ($k=i$). Among them, here we consider the photon-number distribution, the Fano factor, and the noise reduction factor. For all the four increasing values of the pump intensity, our results are consistent with the expected behavior of the TWB. The two arms are well balanced, $\langle m_{s}\rangle\approx\langle m_{i}\rangle$, and each detected-photon distribution is well fitted by the multi-mode thermal distribution in Eq. (1), as shown in Fig. 3(a). The super-Poissonian nature of the measured light is also evidenced by the Fano factor, which is larger than 1 for all the datasets shown in Fig. 3(b). Finally, the noise reduction factor is $R<1$ [see Fig. 3(c)] in all cases, denoting the presence of nonclassical photon-number correlations in the mesoscopic intensity domain. By fitting the latter results using Eq. (4) with $t=1$ (lossless configuration, the two arms are balanced) and $\langle m_{\rm N}\rangle=0$ (noiseless configuration), we can estimate the detection efficiency $\eta=(8.5\pm 0.7)\%$, which is lower than the expected one ($\eta\sim 16\%$ at 693 nm, according to the datasheet[34]). The discrepancy between our estimate and the expected value is likely due to possible losses in the collection of light including a non perfect coupling into fibers. Remarkably, despite its low value, this efficiency is still sufficient to detect nonclassical correlations.
Figure 3: Characterization of the symmetric TWB. (a) Detected-photon distribution $P(m_{k})$ in the idler $k=i$ (blue-scale-colored diamonds) and signal $k=s$ (red-scale-colored circles). The fitting function (line) is the multi-mode thermal distribution (MMTD) in Eq. (1). Light (dark) colors correspond to low (high) $\langle m_{k}\rangle$. (b) Fano factor $F(m_{k})>1$ of each arm as a function of $\langle m_{k}\rangle$. (c) Noise reduction factor $R<1$ as a function of $\langle m_{s}\rangle$. Results in (b) and (c) are obtained by dividing the $10^{5}$ data of each dataset into 4 batches of 25000 shots: the marker with its error bar represents the mean and the associated standard error computed from the batches of the considered dataset.
@@
Figure 4: Characterization of the asymmetric TWB and the noise signals. (a) Detected-photon distribution $P(m_{k})$ in the idler $k=i$ and signal $k=s$ and (b) in the two noise signals (NS). The fitting function (line) is the multi-mode thermal distribution (MMTD) in Eq. (1). (c) Noise reduction factor of the TWB with superimposed noise signal as a function of $\langle m_{\rm N}\rangle$. The dashed line indicates the value of $R$ with associated error in the noiseless lossy configuration. Results in (c) are obtained using the same procedure described in Fig. 3(c).
@@ IV.3 Performance of the discrimination strategy
Figure 5: Effects of the batch size on the estimation of $R$ and $\langle m_{s}\rangle$. Mean value and associated error of the noise reduction factor $R$ (a) and of the mean number of detected photons in the signal arm $\langle m_{s}\rangle$ (b) as a function of the batch size $d_{b}$. Histogram of the values of $R$ (c) and of $\langle m_{s}\rangle$ (d) for $d_{b}=20000$ (top) and $d_{b}=40000$ (bottom). In (c,d) the vertical dashed line denotes the discrimination threshold $R_{th}=0.968$ and $\langle m_{s}\rangle_{th}=3.72$, respectively. Results are obtained from 50005000 batches drawn from the original datasets by bootstrap procedure.
@@
Figure 6: Effects of the batch size and threshold on the discrimination of bits. (a) Error probability (misclassification) as a function of the batch size $d_{b}$ when using $R$ (blue diamonds) or $\langle m_{s}\rangle$ (red circles) to discriminate bits 0 or 1 for a given threshold (see Fig. 5). (b) ROC of each discrimination strategy: True positive rate (TPR) as a function of the false positive rate (FPR) for two representative batch sizes $d_{b}=20000,40000$. The perfect classifier corresponds to (FPR=0,TPR=1) (yellow shadow), while the random classifier to the line TPR=FPR (grey dotted line). The empty markers denote in (a) data associated with $d_{b}=20000,40000$, and in (b) those associated with threshold values $R_{th}=0.968$ and $\langle m_{s}\rangle_{th}=3.72$. Results in (a)-(b) are obtained from 50005000 batches drawn from the original dataset by bootstrap procedure.
@@
Figure 7: Example of discrimination of a 400-bit key for two batch sizes $d_{b}=20000,40000$. The raw key is visualized as a 20×20 binary image, where each pixel corresponds to a single bit. Alice generates her key using a pseudo-random number generator, while Bob obtains his key by measuring either $\langle m_{s}\rangle$ or $R$ and discriminating bits according to the thresholds reported in Fig. 6. Errors in Bob’s keys, highlighted in light- and dark-red, occur with rates determined from 50005000 batches drawn from the original dataset by bootstrap procedure.
@@ V Proof-of-principle test
As expected, Bob can exactly decode Alice’s raw key using the $\langle m_{s}\rangle$-based discrimination strategy, while the raw key he obtains using the $R$-based strategy is affected by errors. In particular, the error probability decreases from $P_{\hbox{\tiny$R$}}^{\rm(err)}\approx 24\%$ to $\approx 16\%$ upon increasing the batch size from $d_{b}=20000$ to $40000$ [see also Fig. 6(a)]. In detail, the false positive/negative rates decreases from ${\rm FPR}\approx 19\%$ to $\approx 11\%$ and ${\rm FNR}\approx 29\%$ to $\approx 22\%$. Therefore, a bit-1 decoded as a bit-0 (i.e., a false negative) constitutes a misclassification error that occurs systematically more often than the opposite case, where a bit-0 is decoded as a bit-1 (i.e., a false positive). This asymmetry in error rates arises from the choice of the discrimination threshold $R_{th}$ in the present example. According to the analysis of the ROC curve, and recalling that ${\rm FNR}\equiv 1-{\rm TPR}$, the condition ${\rm FNR}={\rm FPR}$ corresponds to points lying along the anti-diagonal ${\rm TPR}=1-{\rm FPR}$. However, as illustrated in Fig. 6(b), the points associated with $R_{th}=0.968$ lie below the anti-diagonal, thus indicating ${\rm FNR}>{\rm FPR}$ (see also Fig. 5(c), which shows that the threshold $R_{th}$ lies closer to one distribution than to the other). Since this protocol requires estimating $R$, each bit must be encoded using a batch size at least $d_{b}\sim O(10^{4})$. In this regard, $d_{b}\approx 40000$ represents a suitable batch size to reliably estimate $R$ while keeping the number of shots per bit reasonably limited.
@@
To better investigate the feasibility of the strategy, we simulate an eavesdropping attack by assuming that Eve intercepts part of a batch encoding either 0 or 1, and re-sends a signal with the same mean value as the stolen portion of light. In doing so, Eve replaces a portion of the shots in the signal arm (Bob)—originally nonclassically correlated with those in the idler arm (Alice)—with others from a different light source, which is uncorrelated with Alice’s. This degradation of nonclassical correlations due to the intercept-resend attack is therefore expected to increase the noise reduction factor [see Eq. (4)]. In particular, we consider a synthetic signal generated according to a Poissonian distribution, since this is the statistical distribution that better approximates the distribution of the signal received by Bob in the absence of eavesdropping. The results of this intercept-resend attack are shown in Fig. 8, where the noise reduction factor is shown as a function of the ratio $d_{\rm E}/d_{b}$ between the batch size changed by Eve and that encoding either 0 or 1. The linear dependence of $R$ on $d_{\rm E}/d_{b}$ can be proved analogously to the derivation of Eq. (4). The highlighted bands indicate the uncertainty corresponding to $\pm 2\sigma$. It can be noticed that for bit 0 the value of $R$ exceeds $R^{(0)}+2\sigma$ and $1$ at $d_{\rm E}/d_{b}\approx 9\%$ and $\approx 84\%$, respectively. Instead, for bit 1 it exceeds $R^{(1)}+2\sigma$ and $1$ at $d_{\rm E}/d_{b}\approx 14\%$ and $\approx 78\%$, respectively. This means that Bob can easily detect the presence of Eve by simply checking whether the values of $R$ are larger than $2\sigma$. On the contrary, values $R>1$ are more difficult to observe, at least when Eve, as in the present example, is equipped with an ideal detector that allows perfect estimation of the mean value of the intercepted signal. At the same time, this proof-of-principle test does not take into account possible optimization strategies, such as practical reconciliation or privacy amplification, devised by Bob to improve his receiver.
@@
Figure 8: An ideal intercept-resend attack can be revealed by the values of the noise reduction factor $R$. Eve intercepts $d_{\rm E}$ shots of a batch encoding a bit ($d_{b}=40000$) and resends, in place of the intercepted light, $d_{\rm E}$ shots belonging to a Poissonian light having the same mean value as that of the intercepted light. While the mean value measured by Bob is not affected by this attack (data not shown), the value of the noise reduction factor is. Blue color refers to bit 0, while red color refers to bit 1. The blue dotted line + blue highlighted band corresponds to $R^{(0)}\pm 2\sigma$, while the red dashed line + red highlighted band corresponds to $R^{(1)}\pm 2\sigma$. The values of $R$ corresponding to the intercepted-resent bit 0 (1) are shown as blue open squares (full red squares). Results are averaged over 50005000 realizations of the intercept-resend attack.